OpenStudy (anonymous):

At the end of the first day, 7 weeds appear in your neighborhood park. Each day, the number of weeds increases by four times. How many weeds will be in the park at the end of 14 days?

OpenStudy (stacey):

day 1: 7 weeds day 2: 7 * 4 weeds day 3: (7*4) * 4 = 7 * 4^2 day 4: (7 * 4^2) * 4 = 7 * 4^3 Notice that since we keep multiplying by 4, the exponent on 4 keeps increasing so that on day n, there are [7 * 4^(n-1)] weeds.

OpenStudy (anonymous):

i tried to find the answer but couldnt cuz i dont got a calculaotr so can you answer it please?

OpenStudy (stacey):

469,762,048

OpenStudy (stacey):

If you type 7*4^13 into google, it will also act as a calculator for you.

OpenStudy (anonymous):

626,349,395 392 71 35,456,231 choices ^^^^^^^^

OpenStudy (gw2011):

Stacey's answer is correct. Here is the formula for calculating this problem: (7)(4)^(n-1) where n is the year you are looking for (7)(4)^(14-1)=(7)(4)^13=469,762,048

OpenStudy (gw2011):

Sorry, I wrote incorrectly that n is the year you are looking for, it should n is the day you are looking for.

OpenStudy (stacey):

Okay. So that means that the problem means for the increase to be 4 times the increase of the previous day. day 1: 7 weeds (increase is 7) day 2: 7 + (7 * 4) weeds (increase is 7 * 4) day 3: 7 + (7 * 4) + (7 * 4^2) weeds day 4: 7 + (7 * 4) + (7 * 4^2) + (7 * 4^3) weeds That means we are looking for the sum of a geometric sequence. \[s _{n}=[a*(r ^{n}-1)] / (r-1)\] where a is the first term, r is the ratio (what we multiply by, and n is the number of terms (days).

OpenStudy (stacey):

\[s_{n}=[7∗(4^{14}−1)]/(4−1)\]

OpenStudy (anonymous):

thats not the right answer it has to be one of the choices

OpenStudy (stacey):

If you put 7*(4^14-1)/(4-1) into google, it will calculate it and you will get one of your choices.